graph each line using intercepts
6x-y=3
How do i find my answer

Answer by rfer(16322)

y int x=0
x iny y=0
6x-y=3
6(0)-y=3
y=-3
(0,-3)
---------------
6x-0=3
x=3/6
x=1/2
(1/2,0)